OpenStudy (meowlover17):

The label on the car's antifreeze container claims to protect the car between −40°C and 125°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5 over 9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car. CHECK MY ANSWER???

OpenStudy (meowlover17):

−40 < 5 over 9 (F − 32) < 125; −40 < F < 257

OpenStudy (meowlover17):

−40 > 5 over 9 (F − 32) > 125; −40 > F > 257 −40 < 5 over 9 (F − 32) < 125; −40 < F < 257 −40 < 5 over 9 (F − 32); −40 < F 5 over 9 (F − 32) < 125; F < 257

OpenStudy (meowlover17):

@campbell_st

OpenStudy (meowlover17):

Am i correct?

OpenStudy (campbell_st):

It makes some sense... but its not ideal I'd say you need to convert and get the temps lower \[-40 = \frac{5}{9} (F - 32)~~~so~~~~F = \frac{-40 \times 9}{5} + 32\] same for 125 upper \[125 = \frac{5}{9}(F - 32)~~~or~~~~ F = \frac{125 \times 9}{5}+ 32\] then the inequality is lower < temp < upper

OpenStudy (meowlover17):

So what would my answer be?

OpenStudy (meowlover17):

http://gyazo.com/72bcd794b3821cbc49db6bcf40eef15a Because for this one i got it right using the same type of setup.

OpenStudy (meowlover17):

Hi could you help me?

OpenStudy (meowlover17):

@pooja195 @welshfella

OpenStudy (campbell_st):

well if you do the calculations for the temps in Fahrenheit the the antifreeze temp will be between your 2 calculations.

OpenStudy (meowlover17):

I got it right, i'll medal and fan you since you responded :D thanks.